# BeyondFacial: Identity-Preserving Personalized Generation Beyond Facial Close-ups

### v1

@@ 3 Methods
3.1 Analysis of Facial Close-ups.
@@ 3.2 Dual-Line Inference (DLI)
$\epsilon_{fusion}=\textbf{IdAF}(\epsilon_{id},\epsilon_{semantic})$ (5)
@@ 3.4 Identity Aggregation Prepending (IdAP)
After obtaining the aggregated ID token $\mathcal{T}_{id}^{s}$, we prepend it to the start of semantic tokens $\mathcal{T}_{\text{semantic}}$ (termed ID prepending). Though simple, this step is highly effective: it achieves concatenation (not conflation) of ID and semantic information, preserving the integrity of semantic components. By design, IdAP replaces random identities within $\mathcal{T}_{\text{semantic}}$ with the prepended target ID information, mitigating random ID interference in the final fused identity.
@@ 4.2 Qualitative Evaluation
Gender Expression Enhancement Existing personalized generation methods typically necessitate manual replacement of gender-specific placeholders to align with reference subjects’ gender. For gender-agnostic prompts (e.g., “A person”), generated images often exhibit gender ambiguity. As shown in Fig. 6, our method not only addresses the facial close-up bias but also markedly improves gender expression accuracy for target identities—bolstering overall identity fidelity. This ancillary benefit stems from our IdAP module, which precisely aligns semantic tokens with target identity features, highlighting the framework’s capacity to refine fine-grained identity attributes beyond core identity preservation.
@@ 4.3 Comparison and Analysis
Quantitative Comparison. For quantitative evaluation, we curated a 4,800-sample dataset (12 identities × 20 prompts × 20 seeds) for each method. Since our full-body outputs reduce facial proportions (complicating identity similarity assessment), we used expression-emphasized prompts for identity metrics to ensure facial prominence. For FID, we constructed a 4,800-sample SDXL dataset [23] of full-body, action-rich scenes with random identities—tailored to reflect our focus on generation beyond close-ups.

### v2

@@ 3 Methods
3.1 Analysis of Facial Close-ups
@@ 3.2 Dual-Line Inference (DLI)
$\epsilon_{\text{fusion}}=\textbf{IdAF}(\epsilon_{id},\epsilon_{semantic})$ (5)
@@ 3.4 Identity Aggregation Prepending (IdAP)
After obtaining the aggregated ID token $\mathcal{T}_{id}^{s}$, we prepend it to the start of semantic tokens $\mathcal{T}_{\text{semantic}}$ and get final aggregated semantic token $\mathcal{T}_{semantic}^{id}$ (termed ID prepending). Though simple, this step is highly effective: it achieves concatenation (not conflation) of ID and semantic information, preserving the integrity of semantic components. By design, IdAP replaces random identities within $\mathcal{T}_{semantic}^{id}$ with the prepended target ID information, mitigating random ID interference in the final fused identity.
@@ 4.2 Qualitative Evaluation
Gender Expression Enhancement. Existing personalized generation methods typically necessitate manual replacement of gender-specific placeholders to align with reference subjects’ gender. For gender-agnostic prompts (e.g., “A person”), generated images often exhibit gender ambiguity. As shown in Fig. 6, our method not only addresses the facial close-up bias but also markedly improves gender expression accuracy for target identities—bolstering overall identity fidelity. This ancillary benefit stems from our IdAP module, which precisely aligns semantic tokens with target identity features, highlighting the framework’s capacity to refine fine-grained identity attributes beyond core identity preservation.
@@
Table 3: User Study. Our method demonstrates significant advantages across all four evaluation dimensions.
Identity similarity
Image-text
alignment
Image quality &
visual aesthetics
Scene construction
& figure integrity
AVREGE
InstantID [29] 50.89 48.06 47.20 49.35 48.88
PuLID [8] 57.21 70.95 61.99 68.68 64.71
PhotoMaker [17] 47.95 60.85 56.89 60.50 56.55
IP-Adapter+Photomaker [34] 63.58 48.45 53.35 52.51 54.47
[…]
Figure 8: The visualization of our user study results—presented in the form of a radar chart and a stacked bar chart—clearly illustrates the significant advantages of our method.
@@ 4.3 Comparison and Analysis
Quantitative Comparison.
For quantitative evaluation, we curated a 4,800-sample dataset (12 identities × 20 prompts × 20 seeds) for each method. Since our full-body outputs reduce facial proportions (complicating identity similarity assessment), we used expression-emphasized prompts for identity metrics to ensure facial prominence. For FID, we constructed a 4,800-sample SDXL dataset [23] of full-body, action-rich scenes with random identities—tailored to reflect our focus on generation beyond close-ups.
4.4 User Study
Given the core goal of this work—identity-preserving personalized generation beyond facial close-ups with complete human figures and rich scene semantics—we prioritize not only identity similarity (Face Sim. [4]) but also the integrity of generated humans/scenes and visual aesthetics, distinct from metrics like DINO [37] and CLIP-I [10]. In the main text, we compute FID [11] on a dataset scaled to match SDXL-generated [23] outputs (complete humans, rich scenes), effectively evaluating our method’s substantive progress toward this goal. As shown in Table 1 of the main text, our FID results demonstrate a significant margin over existing works in personalized generation beyond facial close-ups.
To comprehensively assess our method’s advancements, we conduct a user study to gather subjective human evaluations—this study strongly validates practical application potential. Specifically, we set InstantID [29], PhotoMaker [17], PuLID [8], and IP-Adapter+PhotoMaker [34] as baselines, presenting 4 generation tasks (each with reference ID-text-generated image triples for 5 methods: 4 baselines + ours) to each participant. Respondents score 20 combinations (4 tasks × 5 methods) across four criteria using a 100-point scale (minimum increment of 5 points).
  • Identity similarity (generated vs. reference figures): Assesses identity consistency.
  • Image-text alignment (generated content vs. text prompts) : Assesses cross-modal consistency.
  • Image quality & visual aesthetics: Assesses perceptual quality and aesthetic appeal.
  • Scene construction & figure integrity: Assesses completeness of scenes and human figures.
Specifically, we anonymized all method names in the questionnaire and randomly shuffled the method order for each response set. Statistical results are based on feedback from 100 participants, with details presented in Table 3.
The user study results demonstrate our method’s significant advantages over all baselines across the four evaluation criteria (identity similarity, image-text alignment, image quality & visual aesthetics, scene construction & figure integrity), with consistent user approval. This validates the success of our identity-preserving personalized generation (IPPG) beyond facial close-ups. For absolute scores, our method achieves the highest performance in image-text alignment (79.01) and scene construction & figure integrity (78.34). For relative gains, it outperforms the second-ranked method by the largest margins in identity similarity (11.59) and image quality & visual aesthetics (12.39). PuLID [8] ranks second in overall average (64.71) but suffers a notable deficit in identity similarity (57.21), indicating substantial degradation of its ID fidelity with increasing semantic complexity. Equipped with dual ID enhancement, IP-Adapter+PhotoMaker [34] achieves the second-highest identity similarity (63.58)—surpassing other baselines and trailing only ours—yet attains the lowest image-text alignment. This further confirms that balancing ID fidelity and image-text alignment remains a key challenge in contemporary IPPG research.
In contrast to baselines, our method delivers substantial gains across all criteria and achieves a meaningful balance between ID fidelity and cross-modal consistency—addressing the core tradeoff in personalized generation beyond facial close-ups. Fig.8 visualizes the results of our user study in the form of a radar chart and a stacked bar chart, clearly demonstrating the significant advantages of our method.
@@ 4.5 Ablation Study
Figure 11: Application Extension: Consistent action sequence generation. Under the same semantic guidance, BeyondFacial enables action sequence generation for specific characters, incorporating consistency in scenes and costumes.
Figure 12: Application Extension: Multi-scale(zoomable) shot genertion. Under the same semantic guidance, BeyondFacial enables multi-scale shot generation for specific characters—analogous to zoomable lens photography—and facilitates character generation with far-to-near transitions.
@@ 4.6 Applications
4.6.1 Cinematic-Grade Character-Scene generation
To validate our method BeyondFacial’s capability for Cinematic-Grade Character-Scene(CGCS) generation, we first use a large language model (LLM) to generate character-centric story narratives with action and scene transitions. The generated narratives are presented below:
“A person embraces the quiet of early morning along a windswept coastal path, seeking calm amid the rhythm of waves. Dressed in a light gray quick-dry sports jacket, navy blue athletic shorts, white running shoes, and a black silicone sports wristband on the left wrist, their outfit is practical yet sleek—perfect for the crisp sea air. The story unfolds as they move from the path’s start to a cliffside overlook, pausing to absorb the dawn, with salt breezes rustling their hair and golden light gradually warming the horizon.”
Subsequently, we prompt the large language model (LLM) to generate 8 consecutive shot-specific scene descriptions from this narrative, serving as prompts. For a given reference ID, our BeyondFacial method generates 8 character-scene images featuring the reference ID as the core character. Fig.14 and 15 visualize our method’s results for four test IDs. For LLM-generated cinematic shot prompts (complex character-action-scene descriptions), our method achieves personalized generation that balances ID fidelity and semantic expressiveness—delivering complete human figures and visually compelling cinematic scenes. For intuitive comparison, we include results from IP-Adapter+PhotoMaker [34] (Fig.16): its uniform facial close-ups highlight the gap in CGCS generation capability, underscoring our method’s unique value in beyond-facial personalized creation.
Furthermore, we observe from experimental results that BeyondFacial supports consistent action sequence generation (Fig.11) and multi-scale (zoomable) shot generation (Fig.12). This delivers substantial value to IPPG’s application in CGCS generation.
4.6.2 Style and Gender Control Generation
@@
Figure 14: Application Extension: Storyboard Scene Generation Results Based on BeyondFacial (1). Our method achieves personalized generation that balances ID fidelity and semantic expressiveness—delivering complete human figures and visually compelling cinematic scenes.
Figure 15: Application Extension: Storyboard Scene Generation Results Based on BeyondFacial (2). Our method achieves personalized generation that balances ID fidelity and semantic expressiveness—delivering complete human figures and visually compelling cinematic scenes.
Figure 16: Application Extension: Storyboard Scene Generation Results Based on IP-Adapter+Photomaker [34]. IP-Adapter+PhotoMaker’s uniform facial close-ups highlight a notable deficit in character-scene generation.